224 ml of co2 at NTP is passed through  100 ml of 0.2 N solution of NaOH  . Find the weight of Na2Co3 formed .

Grade:12

3 Answers

Viren Sheoran
10 years ago

ANSWER- 0.0106gram

Akash Kumar Dutta
10 years ago

milli equivalents of co2=224000/11200=20
milli equivalents of NaOH=20
hence milli equivants of na2co3 formed=20
mass=20/1000.53=53/50=.16 gm

Akash Kumar Dutta
10 years ago

sorry cal mistake.
ans=.106 gm